Insert the indicated number of arithmetic means given the first and last term. Show your solution.
___1. 3 and 36 indicated number 7

Answer by greenestamps(13200)



7 arithmetic means between a and b means b is equal to a plus 8 times the common difference.

36=3+8d
8d=33
d=33/8 = 4 1/8

3; 7 1/8; 11 1/4; ...

You can do the busy work.